OpenStudy (anonymous):

What is 5.6 × 10−3 yd/year in feet per day? A. 6.813 × 10−7 ft/day B. 6.132 × 10−6 ft/day C. 5 × 10−6 ft/day D. 4.6 × 10−5 ft/day

OpenStudy (anonymous):

5.6x10^-3 yd/year in feet per day. There are 3 feet in a yard (5.6x10^-3)x3=1.68x10^-2 feet/year 365 days in the year \frac{1.68*10^{-2}}{365} = 4.6*10^{-5} the correct answer is D.

OpenStudy (anonymous):

thank you so much